OpenStudy (anonymous):

The density of copper is 559 lb_m/ft^3. What is the mass of the copper pipe in pounds? A copper pipe, used for plumbing from the water main to a house, is 5 meters long and has a circular cross-section 3.2 cm in diameter. There is a 2.4 cm diameter hole in the center of the cross-section for the entire length of the rod.

OpenStudy (mertsj):

pi (3.2/2)^2 * 500 - pi (2.4/2)^2 *500 = 1759.29cm^3

OpenStudy (perl):

now convert to ft^3 , and then multiply by density to get mass

OpenStudy (mertsj):

\[1759.29 cm^3\times\frac{1 IN^3}{2.54^3 cm^3}\times\frac{1 ft^3}{12^3 IN^3}\times\frac{559 lbs}{1 ft^3}\]

OpenStudy (mertsj):

34.73 lbs

OpenStudy (perl):

looks good :)

OpenStudy (perl):

of course the person who pposted it didnt bother waiting for a solution,